OpenStudy (anonymous):

Which investment is better? Investment A: 6% compounded semi annually? Investment B: 4.5% compounded quarterly?

OpenStudy (anonymous):

The answer is $15.13 but I dontk know how to get it

OpenStudy (anonymous):

compare \[\left(1+\frac{0.06}{2}\right)^2\]with \[\left (1+\frac{0.045}{4}\right)^4\] and see which one is larger

OpenStudy (anonymous):

For that solution I got 1.061 for A and 1.046 for B. But the answer that says on the book is $15.13 for A. But I dont know how to get it